Find the measure of an angle and its compliment if one angle measures 24 degrees more than the other
Answer by swincher4391(1107)

By definition if two angles are complementary, they add up to 90 degrees.
Let x be one of the complementary angles. Let x+24 be other.
Then x+x+24 = 90
2x + 24 = 90
2x = 66
x = 33
33 is one angle
(33+24) = 57 is the other.
33,57
-------------------------
To check:
33+57 = 90
Thus, these are complementary angles.